OpenStudy (anonymous):

Write the equation of a circle that has center (2,-4) and passing through (4,-1)

OpenStudy (anonymous):

r^2=(x-2)^2+(y-(-4))^2 hold on i'll find r

OpenStudy (anonymous):

why is it not\[r^2=(x+4)^2+(y-2)^2\]?

OpenStudy (anonymous):

Use distance formula to find r sqrt((delta x)^2+(delta y)^2) delta x just means difference in x which is xf-xi = 4-2 Same for delta y

OpenStudy (anonymous):

oh no wait I remember...

OpenStudy (anonymous):

the formula for a circle with center (h,k) is r^2=(y-k)^2+(x-h)^2

OpenStudy (anonymous):

there is no more info

OpenStudy (anonymous):

yeah, but those links always confuse me and then how do you find r again, isn't there some easy way?

OpenStudy (anonymous):

Use distance formula to find r sqrt((delta x)^2+(delta y)^2) delta x just means difference in x which is xf-xi = 4-2 Same for delta y

OpenStudy (anonymous):

well they tell you the center of the circle, so we know h and k

OpenStudy (anonymous):

yes colorful you have it a little mixed around (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

well whether you write (x-h)^2+(y-k)^2=r^2 or (y-k)^2+(x-h)^2=r^2 it's the same thing, isn't it? and the picture should be something like this